The radius of a circle changes by 1 mm, how will the length of the circle change?
1
Expert's answer
2010-06-08T09:12:55-0400
The length of a circle is expressed by the formula: L = 2*π*R. So if the radius changes by 1 mm (it was R and it became R +/- 1mm), the length of the circle we’ll get: L’ = 2*π*(R +/- 1mm) = 2*π*R +/- 2*π*1mm = L +/- (2*π)mm. So the length of the circle changes by (2*π)mm.